An airforce plan made a trip to dublin and back the trip there took four hours and the trip back took seven hours. it averaged 165 mph faster on the trip there than on the return trip. Find the air force plane's average speed on the out bound trip.
Answer by TimothyLamb(4379)

D/f = 4
D/4 = f
D/r = 7
D/7 = r
f = r + 165
---
D/4 = D/7 + 165
(7/28)D = (4/28)D + 165
(3/28)D = 165
D = 1540 miles
--
Answer:
outbound = 1540/4 = 385 mph
inbound = 1540/7 = 220 mph
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php